OpenStudy (anonymous):

evaluate this integrals; ʃ(√t-1/√t)dt,,,can i ask the solution of the answer???plz....tnx...

OpenStudy (anonymous):

write it as two fraction

OpenStudy (amistre64):

\[\int \frac{\sqrt{t}-1}{\sqrt{t}}\ dt\]??

OpenStudy (anonymous):

\[1 - 1/\sqrt{t}\]

OpenStudy (anonymous):

use exponential notation and then the power rule

OpenStudy (amistre64):

made it to 2000+ now ;)

OpenStudy (anonymous):

\[1-t^{-1\over 2}\]

OpenStudy (anonymous):

ok thank you everyone for answering my question......

OpenStudy (anonymous):

\[\int \sqrt{t}-\frac{1}{\sqrt{t}}dt=\int t^{-\frac{1}{2}}-t^{\frac{1}{2}}dt\]